# A theoretical framework for self-supervised contrastive learning for continuous dependent data

### v3

@@ Abstract
Index Terms:
self-supervised learning, theory, dependent data, time-series
@@ I Introduction
Self-supervised learning (SSL) methods demonstrate significant potential in achieving the quality of supervised models using a vast amount of unlabeled data [1, 2, 3]. The SSL paradigm provides a method for training an encoder using a vast amount of unlabeled data, which produces a universal representation (embedding) of an input signal. Since representations are universal, we can use them to solve various downstream problem tasks without time- and data-consuming full retraining of the encoder [4].
@@
Our key novelties and contributions are the following:
@@ IV-C5 Results
TABLE III: ROC-AUC $(\uparrow)$ values for drought prediction for different forecasting horizons and climatic regions
@@
To evaluate the quality of our embeddings, we use Gradient Boosting as a classifier on top of them to solve the binary classification problem described in Section IV-C1. As an evaluation metric, we use ROC-AUC. The comparison results for different regions are described in Table III. According to the aggregated statistics across all regions, we can see that our method significantly outperforms the standard TS2Vec approach for all forecasting horizons. According to the comparison with the current state-of-the-art SoftCL approach, we observe a slight improvement for all forecasting horizons. Moreover, our similarity matrices are completely interpretable, while this property is not valid for the SoftCL method.
@@ A-A DepTS2Vec for temporal data
Loss function. We use the default instance-wise loss combined with the temporal loss, employing either the MA (4) or AR (5) ground-truth similarity matrices. For AR-based similarity, we chose the best hyperparameter $k$ from the range {1, 5, 10}. For each dataset in the UCR and UEA archives, we identified distinct optimal hyperparameters, including the choice of ground truth similarity matrix and the hyperparameter $k$. As the comprehensive table of optimal hyperparameters is extensive, we have not included it in the text but can provide it upon request.

### v4

@@ I Introduction
Self-supervised learning (SSL) methods demonstrate significant potential in achieving the quality of supervised models using a vast amount of unlabeled data [1, 2, 3, 4]. The SSL paradigm provides a method for training an encoder using a vast amount of unlabeled data, which produces a universal representation (embedding) of an input signal. Since representations are universal, we can use them to solve various downstream problem tasks without time- and data-consuming full retraining of the encoder [5, 6].
@@
Our implementation is publicly available at our GitHub [^1] https://github.com/aayugay99/DepTS2Vec. The key novelties and contributions of this work are as follows:
@@ IV-B5 Results
IV-B6 Ablation study
Table III shows that an optimal similarity measure between samples depends on the temporal characteristics of each dataset type. Specifically, HAR datasets are captured most effectively by short-range dependencies, with AR(k=1) for UCR and MA for UEA, reflecting the short-term nature of human activity signals. In contrast, SENSOR and DEVICE datasets benefit from long-range context, captured effectively by AR(k=10), while ECG signals are best represented with intermediate horizon, AR(k=5). For MOTION, the optimal AR horizon differs across archives: AR(k=5) in UCR and AR(k=1) in UEA. These findings demonstrate that different temporal structures are best captured by different specifications, providing strong empirical support for the flexibility and generality of our framework.
TABLE III: Mean rank $(\downarrow)$ by dataset type and model specification on the UCR and UEA benchmarks.
Type MA AR(k=1) AR(k=5) AR(k=10)
UCR
HAR 3.07 2.00 2.21 2.71
SENSOR 2.82 2.53 2.44 2.21
DEVICE 2.83 2.92 2.29 1.96
ECG 2.29 2.64 2.00 3.07
MOTION 2.58 2.42 2.00 3.00
TRAFFIC 2.25 1.75 3.00 3.00
UEA
HAR 2.11 2.22 2.50 3.17
EEG 2.33 2.50 2.58 2.58
MOTION 3.13 1.13 3.00 2.75
ECG 3.50 3.50 1.50 1.50
Different types of data contain varying amounts of information. Therefore, it is interesting to determine how much information is required to represent a specific type of data. To assess this across all model configurations, we measured the effective dimensionality by counting the number of principal components needed to capture $95\%$ of the variance [41]. We use an embedding size of 320, in accordance with the standard pipeline [16]. As shown in Table IV, at most only $21\%$ of the embedding dimensions are effectively utilized, indicating a dimensional collapse. This is a known property of contrastive learning methods [42], but it also suggests that time series data inherently require fewer latent dimensions for effective representation compared to other domains like computer vision. The stability of this metric implies that the choice of temporal dependency model (hard MA vs. soft AR) impacts downstream task performance more significantly than the expressivity of the embedding space.
TABLE IV: Mean effective dimensionality of embeddings across dataset types and model specifications.
Type MA AR(k=1) AR(k=5) AR(k=10)
UCR
HAR 58.62 58.57 57.71 57.95
SENSOR 65.76 65.82 65.53 65.29
DEVICE 38.42 38.17 38.08 37.92
ECG 28.14 28.57 26.29 26.43
MOTION 54.33 53.83 52.50 52.33
TRAFFIC 3.50 3.50 4.00 4.00
UEA
HAR 55.89 55.67 55.33 55.24
EEG 63.67 63.17 61.67 61.67
MOTION 44.25 44.25 43.75 43.75
ECG 2.00 2.00 2.00 2.00
@@ IV-C5 Results
TABLE V: ROC-AUC $(\uparrow)$ values for drought prediction for different forecasting horizons and climatic regions
@@
To evaluate the quality of our embeddings, we use Gradient Boosting as a classifier on top of them to solve the binary classification problem described in Section IV-C1. As an evaluation metric, we use ROC-AUC. The comparison results for different regions are described in Table V. According to the aggregated statistics across all regions, we can see that our method significantly outperforms the standard TS2Vec approach for all forecasting horizons. According to the comparison with the current state-of-the-art SoftCL approach, we observe a slight improvement for all forecasting horizons. Moreover, our similarity matrices are completely interpretable, while this property is not valid for the SoftCL method.
@@ A-A DepTS2Vec for temporal data
Loss function. We use the default instance-wise loss combined with the temporal loss, employing either the MA (4) or AR (5) ground-truth similarity matrices. For AR-based similarity, we chose the best hyperparameter $k$ from the range {1, 5, 10}. For each dataset in the UCR and UEA archives, we identified distinct optimal hyperparameters, including the choice of ground truth similarity matrix and the hyperparameter $k$. Table III presents the detailed information about the hyperparameters.